37. Jupiter at its present location, were a star? What would Earth be like ? Assume this hypothetical Jupiter with 75 × MJ would be a red dwarf with the same density as the Sun but a surface temperature of only 2500 K. What would be its total luminosity, compared to that of the Sun?

(Hint: recall that to be a star, Jupiter would have to be 75 times more mass than it has today.)
Answer #1


Soin: 02kg = 1.424x1029 kg ST Density of Sun, f= 1.4igm (cm3 = 1410 kg 1 m3 mous og Jupites, mie - 75*my = 75*1-8991021K9 = 1